The capacitors have values C1= 2.0 uf and C2= 4.0 uF, C3=5.0 uF C4= 7.0 uF and the potential difference across the battery is 9.0 V. Assume that the capacitors are connected in series?


1
Expert's answer
2021-04-16T07:25:58-0400

Total capacitance for capacitors in series is CT

"Q=CV"

Were C is Capacitance

V is Voltage supplied

Q IS Charged across the capacitors


"Q_{T}= \frac{1}{c_{1}}+ \frac{1}{c_{2}}+ \frac{1}{c_{3}}+\frac{1}{c_{4}}"


CT= "\frac {1}{2} + \frac {1}{4} + \frac{1}5 + \frac{1}{7} = 0.915uf"


Q= "\ 0.915× 9 = 8.235"


Q= 8.235uC